OpenStudy (aaaa):

Given the function f(x) = 3x + |5 - x|3, find f(8). A. 51 B. 3 C. -3 D. 38

OpenStudy (aaaa):

if u answer i become ur fan

OpenStudy (binary3i):

33

OpenStudy (aaaa):

and u geta a medal

OpenStudy (aaaa):

right answer u get a medal

OpenStudy (aaaa):

nooot

OpenStudy (aaaa):

thats not even an answer choice dude

OpenStudy (ash2326):

remember that |x| = x if x>=0 =-x if x<0 f(x)=3x+|5-x|3 put x=8 f(8)=3*8+|5-8|3 |-3|=3 f(8)=24+3*3 f(8)=33 it's not matching the choices, you'd check the question again

OpenStudy (aaaa):

thamnks to all